### v1

# P1-KAN an effective Kolmogorov Arnold Network for function approximation

Abstract.
A new Kolmogorov-Arnold network (KAN) is proposed to approximate potentially irregular functions in high dimension. We show that it outperforms multilayer perceptrons in terms of accuracy and converges faster. We also compare it with ReLU-KAN, a recently proposed network: it is more time consuming than ReLU-KAN, but more accurate.
1. Introduction
Kolmogorov Arnold Networks [14], based on Arnold Kolmogorov representation theorem, have recently been proposed instead of multilayer perceptrons to approximate functions in high dimension: Arnod and Kolmogorov showed long ago [10] that a multivariate continuous smooth function $f$ on a bounded domain can be written as a finite composition of the sum of continuous functions of a single variable. More precisely, if $f$ is smooth on $[0,1]^{n}$, then
(1) $f(x)=\sum_{i=1}^{2n+1}\psi_{i}(\sum_{j=1}^{n}\Phi_{i,j}(x_{j})),$
where $\Phi_{i,j}:[0,1]\longrightarrow{\mathbb{R}}$ and $\psi_{i}:{\mathbb{R}}\longrightarrow{\mathbb{R}}$.
As the 1D functions can very irregular or even fractal, it has been shown that they may not be learnable in practice [8, 15]. To overcome this limitation, [14] propose to extend this representation. First they propose not to stick to $12n+$ 1 terms in the outer sum in (1) and to define a KAN $l^{th}$ layer as an operator $\psi^{l}_{m,p}$ from $[0,1]^{m}$ to ${\mathbb{R}}^{p}$:
(2) $(\psi^{l}_{m,p}(x))_{k}=\sum_{j=1}^{m}\Phi_{l,k,j}(x_{j}),\mbox{ for }k=1,\ldots,p.$
Second, by stacking the layers, i.e; composing the operator $\psi^{l}$, they define the KAN operator from $[0,1]^{m}$ to ${\mathbb{R}}^{d}$:
(3) $K(x)=(\psi^{L}_{n_{L-1},d}\circ\psi^{L-1}_{n_{L-2},n_{L-1}}\circ\ldots\circ\psi^{1}_{n_{0},n_{1}}\circ\psi^{0}_{m,n_{0}})(x)$
Since all $\psi$ functions are one dimensional, many classical methods are available to propose an easy to implement approximation. In their proposed implementation, [14] use B spline (see for example [4]) associated with the SILU activation function to approximate the $\psi$ function : the spline coefficients and the multiplicative coefficient of the SILU function are learned using a classical stochastic gradient algorithm as done with MLPs.
This network has been rapidly tested replacing MPLs in transformers [24] for example and in various fields : medical sector in [9], vision [5, 12], time series [21]. Strengths and weaknesses of this approach compared to MLPs are discussed in [25] and, depending on its used, its to superiority to MLPs is not always obvious [17, 11].
Following this first article, different evolutions of this architecture are proposed to address different problems :[6] proposes an evolution of the algorithm to replace LSTM in time series, [23] for Graph Collaborative Filtering, [2] for convolutional networks, [1] in mechanics, …
The original spline-based algorithm has several drawbacks. The first disadvantage of this approach is that the spline approximation is very expensive. The second is that the output of a layer may not be in the grid initially chosen for the following layer: the authors propose to adapt the grid during the iterations to the output of the previous layer, which still increases the complexity and the computational cost of the procedure. Finally, since the Kolmogorv representation theorem involves a very irregular function, one may wonder whether it is interesting to use a rather high order approximation as a spline.
To address the first point, many other approximations based on classical numerical analysis have been proposed using: wavelets [3], radial basis [13, 20] which reduces the computation time by 3, Chebyshev polynomials [19] and many others. An interesting representation that leads to a very effective layer is the ReLU-KAN [16] [18], which is based only on the ReLU function, matrix addition and multiplication, and divides the computation time by 20.
To address the second point, some use a sigmoid activation function [19] to get an output in $[0,1]$, others use some adaptation of the support of the basis functions by trying to learn them [16].
Concerned by the possibility of KANs to approximate high dimensional functions in high dimension, especially for stochastic optimisation purposes in [7],[22], we have tested the ReLU-KAN network, sometimes obtaining excellent results in the optimisation of complex hydraulic valleys and sometimes experiencing divergence. In order to avoid this divergence problem observed in operational problems, we have developed the P1-KAN network, borrowing some interesting features from the ReLU-KAN, but clearly defining the support of the layer function and avoiding the network adaptation proposed in [14]. In the first part of the article we describe our architecture. In a second, we compare it with MLPs, ReLU-KAN on function approximation using either regular or very irregular function in different dimensions.
2. The P1-KAN Network
We will first explain the main features of the method, and then go into detail about the algorithm and why P1-KAN is different from other KAN networks.
2.1. The P1-KAN Layer
We assume that the layer is an operator $\kappa$ with support described by $(\underline{x},\bar{x})\in{\mathbb{R}}^{d_{0}}\times{\mathbb{R}}^{d_{0}}$ and with values in ${\mathbb{R}}^{d_{1}}$. As for the Relu-KAN layer, a number of meshes per direction $M$ are used to discretize $[\underline{x}^{i}_{1}\bar{x}^{i}_{1}]\times\ldots\times[\underline{x}^{i}_{d_{1}}\bar{x}^{i}_{d_{1}}]$, giving the mesh vertices $(\hat{x}^{i}_{j})_{1\leq j\leq M-1}$ of trainable variables in $]\underline{x}^{i},\bar{x}^{i}[$. Set $\hat{x}^{i}_{0}=0$, $\hat{x}^{i}_{M}=1$, the function $\Phi^{i}$ in (2) is discretized using a P1 finite element method: for $x\in[\underline{x}^{i},\bar{x}^{i}]$,
(4) $\Phi^{i}(x)=\sum_{j=0}^{M}a^{i}_{j}\Psi^{i}_{j}(x)$
where $(a^{i}_{j})_{j=0,M}$ are also trainable variables and $(\Psi^{i}_{j})_{j=0,M}$ is the basis of the shape function $\Psi^{i}_{j}$ with compact support in each interval $[\hat{x}^{i}_{j-1},\hat{x}^{i}_{j+1}]$ for $j=1,\dot{M}-1$ and defined as:
(7) $\Psi^{i}_{j}(x)=\left\{\begin{array}[]{cc}\frac{x-\hat{x}^{i}_{j-1}}{\hat{x}^{i}_{j}-\hat{x}^{i}_{j-1}}&x\in[\hat{x}^{i}_{j-1},\hat{x}^{i}_{j}]\\ \frac{\hat{x}^{i}_{j+1}-x}{\hat{x}^{i}_{j+1}-\hat{x}^{i}_{j}}&x\in[\hat{x}^{i}_{j},\hat{x}^{i}_{j+1}]\\ \end{array}\right.\mbox{ such that }\Psi^{i}_{j}(\hat{x}^{i}_{k})=\delta_{k,j},$
for $j=1,\ldots,M-1$. Similarly, $\Psi^{i}_{0}$ (or $\Psi^{i}_{M}$) is defined as a linear by-part function with support $[\underline{x}^{i},\hat{x}^{i}_{1}]$ (or $[\hat{x}^{i}_{M-1},\bar{x}^{i}]$) and such that $\Psi^{i}_{0}(\underline{x}^{i})=1$ (or $\Psi^{i}_{M}(\bar{x}^{i})=1$). Unlike other networks, the P1-KAN layer, which is theoretically described as an operator from ${\mathbb{R}}^{d_{0}}$ to ${\mathbb{R}}^{d_{1}}$ by the equation (2), takes as input not only a sample $x\in{\mathbb{R}}^{d_{0}}$ but also the description of the support $(\underline{x}^{i}\bar{x}^{i})$.
In detail, the vertices in $]\underline{x}^{i}\bar{x}^{i}[$ are generated for each direction $i$ for $1\leq j<M$ by
(8) $\hat{x}_{j}^{i}=\underline{x}^{i}+(\bar{x}^{i}-\underline{x}^{i})\frac{\sum_{k=1}^{j}e^{-y_{k,i}}}{\sum_{k=1}^{M}e^{-y_{k,i}}}$
where the matrix $Y=(y_{k,i})_{1\leq k\leq M,1\leq i\leq d_{0}}$ has elements in ${\mathbb{R}}$. The operator value is given by:
(9) $\kappa(x)_{k}=\sum_{i=1}^{d_{0}}\sum_{j=0}^{M}a_{k,j,i}\Psi^{i}_{j}(x_{i}),\mbox{ for }k=1,\ldots,d_{1}.$
The tensor $A=(a_{k,j,i})_{0\leq j\leq M,1\leq i\leq d_{0},1\leq k\leq d_{1}}$ and $Y$ are the trainable variables of the network.
As output, the layer returns the values of $\kappa(x)$ in ${\mathbb{R}}^{d_{1}}$ and the lattice $G=[\underline{G},\bar{G}]$ obtained from the possible $\kappa(x)$ values. Due to the use of the $P_{1}$ finite element approximation, this output lattice is exactly obtained from the $A$ tensor by:
$\underline{G}_{k}=$ $\sum_{i=1}^{d_{0}}\min_{0\leq j\leq M}a_{k,j,i}$
$\bar{G}_{k}=$ $\sum_{i=1}^{d_{0}}\max_{0\leq j\leq M}a_{k,j,i}$
for $1\leq k\leq d_{1}$.
2.2. The global P1-KAN network
As shown in the previous section, the P1-KAN layer inputs $x$ values and a grid, and it outputs the values obtained by the operator and a grid. Therefore, it is natural to stack the layers without using any grid adaptation or using any sigmoid function to send the output of the layer back to a known bounded domain.
The P1-KAN network takes for the initial grid used for the first layer a grid corresponding to the bounded domain where we want to approximate the unknown function. An implementation in Tensorflow is available at https://fime-lab.org/warin-xavier.
3. Numerical results
In this section we compare the classical feedforward network with ReLU-KAN and P1-KAN on two types of functions defined on $[0,1]^{d}$.
  1. A
    The first function is regular but very fast oscillating with increasing dimension and is defined for $x\in[0,1]^{d}$
    (10) $f(x)=\cos(\sum_{i=1}^{d}iy_{i})$
    where $y=0.5+\frac{2x-1}{\sqrt{d}}$. The function in 2D is given in figure 1.
    Figure 1. function A in 2D
  2. B
    The second is a very irregular one and is given as
    (11) $f(x)=d(\prod_{i=1}^{d}y_{i}+2(4\prod_{i=1}^{d}x_{i}-\lfloor 4\prod_{i=1}^{d}x_{i}\rfloor)-1)$
[…]
    Figure 2. function B in 2D
To approximate a function $f$ with a neural network $\kappa$ we use the classical quadratic loss function defined as:
$L={\mathbb{E}}[(f(X)-\kappa(X))^{2}]$
and where $X$ is a uniform random variable on $[0,1]^{d}$. Using Tensorflow, a stochastic gradient algorithm with the ADAM optimizer, a learning rate of $31e-$ 3, and a batch size of $1000$, we minimize the loss $L$. The MLPs use a ReLU activation function, using either 2 layers with 10, 20 ,40 neurons for each layer or 3 layers with 10, 20, 40, 80, 160 neurons: In each case, the MLP are optimized by varying the number of neurons and layers, and only the result that gives the smallest loss during the iterations is kept for the plots. The two KAN networks are compared using the same parameterization (number of hidden layers, number of neurons, number of meshes used for the 1D functions). The ReLU-KAN has an additional parameter $k$, which we keep at $3$ as suggested in the original article. For all plots, every 100 gradient iterations, the loss is calculated more accurately using $10^{5}$ samples, giving a series of log-losses plotted using a moving average window of 10 results.
ReLU-KAN is very efficient in terms of computation time as it can be broken down into a few operations involving only the ReLU function, matrix addition and multiplication. On a 11th generation Intel(R) Core(TM) i7-11850H @ 2.50GHz, using the same parameterisation of the KAN nets, the P1-KAN computation time is between 1.5 and 2 times slower than the ReLu-KAN.
3.1. Results for the A function
The results in dimension 6 shown in figure 3 indicate that the P1-KAN network converges faster than the other networks if $M$ is high enough. In general, the ReLU-KAN network converges at least as well as the best feedforward.
2 hidden layers of 10 neurons, $M=5$
3 hidden layers of 10 neurons, nMesh$=5$
2 hidden layers of 10 neurons, $M=10$
3 hidden layers of 10 neurons, nMesh$=10$
2 hidden layers of 10 neurons, $M=20$
3 hidden layers of 10 neurons, $M=20$
Figure 3. Results in dimension 6 for function A
In dimension 12 on figure 4, P1-KAN again gives the best results when $M$ is small, but doesn’t give any results when $M$ is too large. The Relu-KAN network always fails, while the feedforward network seems to converge very slowly, but the accuracy remains limited.
2 hidden layers of 10 neurons, $M=5$
3 hidden layers of 10 neurons, $M=5$
3 hidden layers of 10 neurons, $M=10$
3 hidden layers of 10 neurons, $M=10$
Figure 4. Results in dimension 12 for function A
In dimension 13 on figure 5, we give the only plot where the networks have converged. Only the P1-KAN network succeeds by using a small $M$ in high dimension.
3 hidden layers of 10 neurons, $M=5$
3 hidden layers of 10 neurons, $M=5$
Figure 5. Results in dimension 13 for function A
Above dimension 13, all methods have failed. These results indicate that for complex smooth functions, the P1-KAN network is the most effective and that it is optimal to keep $M$ small.
3.2. Results for the B function
In dimension 2 on figure 6, we see that the feedforward lags far behind the KAN networks, both of which perform well for small $M$ values. For higher values of $M$, P1-KAN is the only network that gives very good results.
2 hidden layers of 10 neurons, $M=5$
3 hidden layers of 10 neurons, $=5$
2 hidden layers of 10 neurons, $M=10$
3 hidden layers of 10 neurons, nMesh$=10$
2 hidden layers of 10 neurons, $M=20$
3 hidden layers of 10 neurons, $m=20$
Figure 6. Results in dimension 2 for function B
The previous results are confirmed in dimension 3 in figure 6, but we can see that in some configurations the ReLU-Kan network can face some converge problems.
2 hidden layers of 10 neurons, $M=10$
3 hidden layers of 10 neurons, nMesh$=10$
2 hidden layers of 10 neurons, $M=20$
3 hidden layers of 10 neurons, $m=20$
Figure 7. Results in dimension 3 for function B
Finally, going up to dimension 5, we see that the ReLU-KAN network can diverge. The P1-KAN network is the only one that gives acceptable results.
2 hidden layers of 10 neurons, $M=10$
3 hidden layers of 10 neurons, nMesh$=10$
2 hidden layers of 10 neurons, $M=20$
3 hidden layers of 10 neurons, $m=20$
Figure 8. Results in dimension 5 for function B
4. Conclusion
The P1-KAN network appears to be an excellent network for approximating functions. We have been able to reproduce on simple examples the divergence problems encountered using the ReLU-KAN on operational problems when optimising complex hydraulic valleys. Further comparisons using MLPs and KANs to optimize French hydraulic valleys will be presented in a forthcoming article.